# Macro Graph of Experts for Billion-Scale Multi-Task Recommendation

### v1

@@ B.1. Dataset Details
B.2. Baseline Details
@@
We compare our proposed MGOE with ten relevant representative multi-task models, which can be roughly divided into two categories.
(I) Traditional Multi-task Models:
  • MMoE (Ma et al., 2018b) introduces a gating network to assign different expert networks to different tasks, allowing the model to share knowledge between different tasks while retaining the task-specific information processing ability.
  • PLE (Tang et al., 2020) extracts deep information layer by separating shared experts and task specific experts, which helps to solve the negative transfer problem and seesaw problem in multi-task learning.
  • ESMM (Ma et al., 2018a) effectively solves the problem of data sparsity and sample selection bias in CVR estimation, and improves the generalization ability of the model by training on the entire sample space.
  • AITM (Xi et al., 2021) proposes an adaptive information transfer multitask framework to models the sequence dependencies between user multi-step transitions through an adaptive information transfer module.
  • STEM (Su et al., 2024) proposes a novel share and task-specific embedding paradigm and it is equipped with an all forward task-specific backward gating network to facilitate the learning of task-specific embeddings and direct knowledge transfer across tasks.
  • MoME (Xu et al., 2024) designs an expert to share parameter efficiently and achieve resource utilization. It extracts from an identical over-parameterized base network by utilizing a binary mask learning mechanism composed of neuron-level model masking and weight-level expert masking.
(II) Graph-based Methods:
  • MoGENet (Hu et al., 2022) introduce a multi-channel graph neural network module designed to adaptively extract the most pertinent and significant information from the bipartite graph.
  • MMoCEG (Zhang et al., 2023b) addresses the expert representation degeneration problem through contrastive learning, and introduces a graph-based gating mechanism to recognize the commonalities and differences among multiple regions.
  • POGCN (Zhang et al., 2024a) introduces partial order graphs to unify separate behavior graphs into a joint graph, leveraging graph convolutional networks to aggregate neighbor information while considering behavior relations.
  • MacGNN (Chen et al., 2024) resolves the computational complexity problems in the infrastructure by reducing the node count from billions to hundreds, it improves the accuracy of CTR prediction while remaining computationally efficient.

### v4

@@ Appendix A Experimental Details
Table 6. Multi-task recommendation performance comparison results over five trial runs.
Taobao Click AUC 0.5577±0.0005 0.5573±0.0006 0.5482±0.0026 0.5575±0.0002 0.5549±0.0009 0.5719±0.0040 0.5793±0.0138 0.5725±0.0049 0.5677±0.0049 0.6039±0.0009 0.6125±0.0035 1.42%
Logloss 0.3744±0.0017 0.3876±0.0016 0.3715±0.0003 0.3781±0.0032 0.3930±0.0105 0.4376±0.0176 0.4245±0.0273 0.4475±0.0791 0.3688±0.0004 0.3678±0.0011 0.3632±0.0015 1.25%
Favor AUC 0.5665±0.0013 0.6140±0.0011 0.5760±0.0099 0.5668±0.0011 0.6126±0.0026 0.6493±0.0106 0.6057±0.0365 0.6178±0.0125 0.5485±0.0108 0.6282±0.0048 0.7961±0.0033 22.61%
Logloss 0.1527±0.0006 0.1577±0.0025 0.1493±0.0004 0.1548±0.0016 0.1614±0.0036 0.1695±0.0074 0.1771±0.0094 0.1712±0.0239 0.1575±0.0038 0.1494±0.0008 0.1321±0.0015 11.52%
Cart AUC 0.5460±0.0008 0.5692±0.0010 0.5408±0.0025 0.5451±0.0008 0.5668±0.0016 0.5745±0.0047 0.5724±0.0177 0.5603±0.0044 0.5549±0.0031 0.5935±0.0026 0.6647±0.0014 12.00%
Logloss 0.2479±0.0007 0.2589±0.0018 0.2456±0.0005 0.2533±0.0028 0.2627±0.0072 0.2899±0.0054 0.2949±0.0153 0.2962±0.0501 0.2485±0.0005 0.2432±0.0007 0.2359±0.0004 3.00%
Buy AUC 0.5542±0.0014 0.5662±0.0041 0.5143±0.0027 0.5576±0.0039 0.5686±0.0010 0.5795±0.0061 0.6414±0.0135 0.5989±0.0045 0.6057±0.0033 0.6869±0.0069 0.7012±0.0077 2.08%
Logloss 0.0968±0.0001 0.0993±0.0017 0.0970±0.0002 0.0995±0.0014 0.1096±0.0039 0.1132±0.0042 0.1140±0.0047 0.1075±0.0104 0.1020±0.0018 0.0922±0.0002 0.0915±0.0005 0.76%
Average AUC 0.5561±0.0005 0.5767±0.0014 0.5448±0.0026 0.5567±0.0011 0.5757±0.0005 0.5938±0.0063 0.5997±0.0203 0.5874±0.0066 0.5692±0.0055 0.6281±0.0038 0.6936±0.0040 10.43%
Logloss 0.2179±0.0008 0.2259±0.0018 0.2159±0.0002 0.2214±0.0021 0.2317±0.0054 0.2526±0.0086 0.2526±0.0141 0.2556±0.0409 0.2192±0.0016 0.2119±0.0007 0.2057±0.0010 2.93%
QB-video Click AUC 0.7876±0.0031 0.7935±0.0009 0.7949±0.0002 0.7949±0.0003 0.7925±0.0008 0.8240±0.0044 0.6425±0.1038 0.8219±0.0220 0.8350±0.0035 0.8323±0.0040 0.8497±0.0070 1.76%
Logloss 0.4592±0.0021 0.4564±0.0008 0.4537±0.0003 0.4549±0.0011 0.4577±0.0032 0.4400±0.0045 3.7794±3.9909 0.4348±0.0176 0.4222±0.0035 0.4402±0.0105 0.4037±0.0067 4.38%
Like AUC 0.6385±0.0037 0.6279±0.0039 0.6251±0.0075 0.7518±0.0113 0.7593±0.0054 0.7396±0.0112 0.5503±0.0343 0.7396±0.0302 0.7264±0.0076 0.6953±0.0143 0.8431±0.0090 11.04%
Logloss 0.0406±0.0001 0.0380±0.0010 0.0367±0.0001 0.0361±0.0010 0.0388±0.0022 0.0342±0.0002 0.1509±0.0810 0.0359±0.0016 0.0379±0.0005 0.0404±0.0008 0.0310±0.0012 9.36%
Share AUC 0.5105±0.0041 0.5296±0.0180 0.5099±0.0017 0.5523±0.0080 0.5281±0.0562 0.5529±0.0345 0.5444±0.0095 0.5809±0.0382 0.5347±0.0075 0.5866±0.0230 0.6630±0.0189 13.02%
Logloss 0.0088±0.0001 0.0083±0.0002 0.0086±0.0001 0.0084±0.0001 0.0156±0.0023 0.0082±0.0001 0.0327±0.0055 0.0196±0.0140 0.0112±0.0005 0.0081±0.0001 0.0080±0.0001 1.23%
Follow AUC 0.6037±0.0035 0.6057±0.0069 0.5470±0.0070 0.6137±0.0035 0.6031±0.0031 0.5932±0.0368 0.5796±0.0476 0.6579±0.0200 0.6678±0.0053 0.6560±0.0211 0.7301±0.0107 9.33%
Logloss 0.0068±0.0001 0.0066±0.0001 0.0067±0.0002 0.0063±0.0001 0.0065±0.0001 0.0073±0.0002 0.0315±0.0062 0.0121±0.0093 0.0100±0.0004 0.0072±0.0001 0.0060±0.0001 4.76%
Average AUC 0.6351±0.0013 0.6392±0.0054 0.6145±0.0038 0.6781±0.0057 0.6707±0.0361 0.6774±0.0217 0.5792±0.0488 0.7001±0.0276 0.6910±0.0060 0.6926±0.0156 0.7715±0.0114 10.20%
Logloss 0.1289±0.0006 0.1273±0.0004 0.1264±0.0001 0.1264±0.0005 0.1296±0.0019 0.1224±0.0012 0.9986±1.0209 0.1256±0.0106 0.1203±0.0012 0.1240±0.0028 0.1124±0.0020 6.57%
QK-article Read AUC 0.6996±0.0008 0.7080±0.0020 0.7135±0.0034 0.7105±0.0047 0.7256±0.0006 0.7451±0.0045 0.7613±0.0073 0.7589±0.0019 0.7352±0.0028 0.7741±0.0049 0.7937±0.0078 2.53%
Logloss 0.1185±0.0013 0.1487±0.0028 0.1161±0.0005 0.1291±0.0032 0.1186±0.0029 0.1287±0.0081 0.1309±0.0182 0.1149±0.0037 0.1145±0.0024 0.1047±0.0088 0.0921±0.0060 12.03%
Like AUC 0.7768±0.0016 0.7947±0.0022 0.8107±0.0059 0.8003±0.0042 0.8184±0.0013 0.7892±0.0085 0.7322±0.0276 0.8284±0.0067 0.7763±0.0267 0.8670±0.0039 0.9250±0.0074 6.69%
Logloss 0.0782±0.0005 0.0933±0.0025 0.0722±0.0003 0.0845±0.0025 0.0741±0.0011 0.0776±0.0045 0.0905±0.0053 0.0770±0.0043 0.0811±0.0021 0.0698±0.0033 0.0532±0.0023 23.78%
Favor AUC 0.8001±0.0015 0.8231±0.0043 0.8303±0.0096 0.8227±0.0038 0.8287±0.0026 0.7996±0.0115 0.7447±0.0310 0.8480±0.0061 0.7760±0.0264 0.8835±0.0047 0.9351±0.0070 5.84%
Logloss 0.0369±0.0001 0.0402±0.0007 0.0325±0.0002 0.0395±0.0012 0.0341±0.0008 0.0343±0.0005 0.0468±0.0038 0.0354±0.0015 0.0409±0.0021 0.0342±0.0018 0.0251±0.0004 22.77%
Follow AUC 0.7040±0.0050 0.7385±0.0028 0.6936±0.0151 0.7421±0.0044 0.7304±0.0059 0.7150±0.0100 0.6295±0.0524 0.7597±0.0093 0.7072±0.0165 0.7863±0.0044 0.8225±0.0085 4.60%
Logloss 0.0078±0.0001 0.0088±0.0001 0.0076±0.0001 0.0089±0.0002 0.0074±0.0001 0.0076±0.0002 0.0193±0.0056 0.0077±0.0002 0.0154±0.0027 0.0075±0.0002 0.0066±0.0001 10.81%
Average AUC 0.7451±0.0014 0.7661±0.0015 0.7620±0.0024 0.7689±0.0038 0.7758±0.0014 0.7622±0.0086 0.7169±0.0296 0.7988±0.0060 0.7487±0.0181 0.8277±0.0045 0.8691±0.0070 5.00%
Logloss 0.0603±0.0005 0.0728±0.0011 0.0571±0.0002 0.0655±0.0002 0.0586±0.0011 0.0620±0.0033 0.0719±0.0082 0.0588±0.0024 0.0630±0.0023 0.0540±0.0035 0.0442±0.0022 18.15%
A.1. Dataset Details
  • Taobao [^2] https://tianchi.aliyun.com/dataset/649 (Zhu et al., 2018) is a user behavior dataset provided by Alibaba, one of the most famous online shopping malls in China. It includes the activities between 28,910 users and 65,087 items from November to December 2017, including clicks, favors, carts, and buys.
  • QB-video [^3] https://github.com/yuangh-x/2022-NIPS-Tenrec/tree/master (Yuan et al., 2022) is a large-scale real-world dataset obtained from the recommendation data of two content platform applications of Tencent. We select a subset of video scenes that consists of 28,803 users and 15,540 items, including four basic behaviors: click, like, share, and follow.
  • QK-article33footnotemark: 3 (Yuan et al., 2022) is a content dataset under the article scene provided by Tencent. It records the behaviors of 52,052 users and 30,273 products, which mainly fall into four categories: reading, liking, favoring, and following.
A.2. Baseline Details
@@
We compare our proposed MGOE with thirteen relevant representative models, which can be roughly divided into three categories.
Traditional Multi-Task Models: (i) MMoE (Ma et al., 2018a) introduces a gating network to assign different expert networks to different tasks. (ii) PLE (Tang et al., 2020) extracts deep information layer by separating shared experts and task specific experts. (iii) ESMM (Ma et al., 2018b) solves the problem of data sparsity and sample selection bias and improves generalizability by training the entire sample space. (iv) AITM (Xi et al., 2021) proposes an adaptive information transfer framework to model sequence dependencies. (v) STEM (Su et al., 2024) proposes a share and task-specific embedding paradigm to facilitate the learning of task-specific embeddings and direct knowledge transfer across tasks. (vi) MoME (Xu et al., 2024) designs an expert to share parameters efficiently and maximize resource utilization.
Graph-based Methods: (i) MoGENet (Hu et al., 2022) introduces a multi-channel GNN module to extract information from the bipartite graph. (ii) MMoCEG (Zhang et al., 2023a) introduces a graph-based gating mechanism to recognize commonalities and differences between multiple regions. (iii) POGCN (Zhang et al., 2024) introduces partial order graphs utilizing GNN to consider behavior relations. (iv) MacGNN (Chen et al., 2024) solves computational complexity problems in the infrastructure by reducing the number of nodes from billions to hundreds.
Single-Task Learning Methods: (i) DeepFM (Guo et al., 2017) combines a factorization machine and a DNN to model user-item relationships. (ii) DIN (Zhou et al., 2018) introduces an attention mechanism to capture the correlation between user’s historical behavior and candidate ads. (iii) DIEN (Zhou et al., 2019) further improves the accuracy of the CTR prediction by capturing the evolution of the user interests.
A.3. Experimental Results with Five Runs
This subsection provides detailed results of experiments where all models were run five times independently on each dataset. The reported metrics are presented as mean ± standard deviation, ensuring statistical robustness and reliability of the performance comparisons. From Table 6, we observe that MGOE also attains the highest average AUC values and the lowest average Logloss values across all task types. Specifically, in terms of the AUC metric, MGOE surpasses the best baseline with average AUC improvements of 10.43% on Taobao, 10.2% on QB-video, and 5% on QK-article. Regarding the Logloss metric, MGOE achieves average effective gains of 2.93%, 6.57%, and 18.15% respectively. These results further validate the consistency and stability of MGOE superiority across different runs.
A.4. Performance Comparison with STL Models
To further validate the effectiveness of MGOE, we conducted experiments using three STL models (DeepFM, DIN, and DIEN) across the three datasets. The results, as presented in Table 7, demonstrate that our model outperforms these STL baselines in both AUC and Logloss metrics, reinforcing its superiority.
Table 7. Comparison results with the STL models.
Models DeepFM DIN DIEN MGOE Improv.(%)
Taobao Click AUC 0.5501±0.0252 0.5632±0.0004 0.5644±0.0003 0.6125±0.0035 8.52%
Logloss 0.6280±0.0695 0.3743±0.0010 0.3753±0.0011 0.3632±0.0015 2.97%
Favor AUC 0.5810±0.0494 0.5721±0.0006 0.5861±0.0004 0.7961±0.0033 35.83%
Logloss 0.2235±0.0218 0.1527±0.0006 0.1523±0.0001 0.1321±0.0015 13.26%
Cart AUC 0.5432±0.0218 0.5660±0.0009 0.5670±0.0010 0.6647±0.0014 17.23%
Logloss 0.3705±0.0511 0.2463±0.0007 0.2462±0.0003 0.2359±0.0004 4.18%
Buy AUC 0.5655±0.0291 0.6191±0.0014 0.6187±0.0007 0.7012±0.0077 13.26%
Logloss 0.1401±0.0163 0.0942±0.0008 0.0946±0.0006 0.0915±0.0005 2.87%
Average AUC 0.5600±0.0314 0.5801±0.0005 0.5840±0.0006 0.6936±0.0040 18.77%
Logloss 0.3405±0.0397 0.2168±0.0006 0.2171±0.0005 0.2057±0.0010 5.12%
QB-video Click AUC 0.6948±0.1238 0.8010±0.0029 0.8013±0.0014 0.8497±0.0070 6.04%
Logloss 0.7279±0.4617 0.4419±0.0030 0.4427±0.0021 0.4037±0.0067 8.64%
Like AUC 0.6177±0.1341 0.6631±0.0026 0.6640±0.0021 0.8431±0.0090 26.97%
Logloss 0.1633±0.1037 0.0406±0.0001 0.0406±0.0001 0.0310±0.0012 23.65%
Share AUC 0.5651±0.0443 0.5571±0.0032 0.5612±0.0027 0.6630±0.0189 17.32%
Logloss 0.0267±0.0144 0.0084±0.0006 0.0081±0.0008 0.0080±0.0001 1.23%
Follow AUC 0.6169±0.0691 0.6130±0.0033 0.6271±0.0022 0.7301±0.0107 16.42%
Logloss 0.0234±0.0125 0.0079±0.0004 0.0074±0.0006 0.0060±0.0001 18.92%
Average AUC 0.6236±0.0928 0.6585±0.0030 0.6634±0.0021 0.7715±0.0114 16.29%
Logloss 0.2353±0.1481 0.1247±0.0011 0.1247±0.0009 0.1124±0.0020 9.86%
QK-article Read AUC 0.6013±0.1241 0.7817±0.0044 0.7810±0.0045 0.7937±0.0078 1.54%
Logloss 0.7391±0.5120 0.1012±0.0061 0.1022±0.0059 0.0921±0.0060 8.99%
Like AUC 0.6157±0.1505 0.8583±0.0080 0.8586±0.0075 0.9250±0.0074 7.73%
Logloss 0.4124±0.2750 0.0668±0.0023 0.0662±0.0021 0.0532±0.0023 19.64%
Favor AUC 0.6198±0.1447 0.8571±0.0087 0.8610±0.0082 0.9351±0.0070 8.61%
Logloss 0.1655±0.1061 0.0331±0.0021 0.0327±0.0020 0.0251±0.0004 23.24%
Follow AUC 0.6046±0.1062 0.7765±0.0105 0.7777±0.0100 0.8225±0.0085 5.76%
Logloss 0.0252±0.0138 0.0073±0.0087 0.0074±0.0085 0.0066±0.0001 9.59%
Average AUC 0.6104±0.1314 0.8184±0.0079 0.8195±0.0076 0.8691±0.0070 6.05%
Logloss 0.3356±0.2267 0.0521±0.0048 0.0521±0.0046 0.0442±0.0022 15.16%
A.5. Complexity Analysis
We further compare the complexity with the graph-based and the expert-based methods to further validate the efficiency of MGOE.
For the graph-based methods, for example, MoGENet, given the number of nodes $m+n$, the number of layers $L$ and the feature dimension $d$, the approximate complexity for feature passing in one layer is $O((m+n)d^{2})$ and the node feature transformation is $O((m+n)d)$. The overall complexity for $L$ layers is $O(L(m+n)(1+d)d)$. When $m+n$ become extremely large, this complexity significantly raised. For the expert-based methods (e.g., PLE and STEM), let $h_{e}$, $h_{s}$, $h_{g}$, and $h_{o}$ be the number of shared expert neurons, task-specific expert neurons, gate networks neurons, and tower neurons respectively. The complexity of expert layer is approximately $O(d(h_{e}+h_{g}+h_{s}))$. For a $L$ expert layers method, the complexity is $O(Ld(h_{e}+h_{g}+h_{s})+h_{e}h_{o})$ as the number of tasks/shared expert/task-specific expert is a constant. For MGOE, let $\widetilde{m}+\widetilde{n}$ be the number of macro nodes, the complexity of each macro graph expert is $O((\widetilde{m}+\widetilde{n})^{2}d)$. Since there are $T+1$ modules, the total complexity is $O((T+1)(\widetilde{m}+\widetilde{n})^{2}d)$. The macro and micro fusion are composed of CGC layers so the complexity is also $O(Ld(h_{e}+h_{g}+h_{s}))$. The overall complexities of MGOE is approximately $O((\widetilde{m}+\widetilde{n})^{2}d+Ld(h_{e}+h_{g}+h_{s})+h_{e}h_{o})$. Since the number of macro nodes $\widetilde{m}+\widetilde{n}$ is relatively small, the complexity of MGOE is comparable to that of the expert-based methods.
Overall, MGOE strikes a better balance between computational efficiency and representational power compared to them.
A.6. Pseudocode
The complete procedures of MGOE are outlined in Algorithm 1.
Input : Task set $\mathcal{T}$; Task-specific dataset $\mathcal{D}_{t}$
Output : Predicted labels
Macro Task Merging Graph
Merge interaction matrices into $\mathbf{S}$} using Equation 7;
Construct macro nodes and macro edges by Equation 10 and Equation 11;
Build macro task table $T$;
Macro Graph of Experts
Initialize model parameters $\Theta$;
for each mini-batch do
   Generate micro user/item embeddings ${\bm{e}}_{u}$, ${\bm{e}}_{i}$ and macro embeddings $\widetilde{{\mathbf{E}}}(\mathcal{\widetilde{N}}_{u})$, $\widetilde{{\mathbf{E}}}(\mathcal{\widetilde{N}}_{i})$;
 foreach $t\in\mathcal{T}$ do
      Compute macro weights $w$ and aggregated macro embeddings $\widetilde{{\mathbf{Z}}}$ via Equation 15 and Equation 16;
      Calculate macro graph expert output $\operatorname{MGE}_{t}$ by Equation 18;
      Compute micro neighbor representations $e_{\text{nbr}}^{t}$ using Equation 19;
      Concatenate micro and macro representations and pass through expert layer $h_{t}$ by Equation 20 and Equation 21;
   end foreach
  Compute final prediction $\hat{\mathrm{Y}}_{ui}$ by Equation 22 and Equation 23;
   Update model parameters $\Theta$ via $\mathcal{L}_{\text{MGOE}}$;
end for
return Predicted labels;
Algorithm 1 MGOE